OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (sidsiddhartha):

\[f(n)=f(1)+f(2)+f(n-1)\] now here put \(n\)=3 u'll get value of f(3)

OpenStudy (sidsiddhartha):

after that put n=4

OpenStudy (anonymous):

f(3)=f(1)+f(2)+f(n-1)=2+4+2=8

OpenStudy (anonymous):

2+4+3=9

OpenStudy (sidsiddhartha):

what is that?

OpenStudy (anonymous):

f(4)=f(1)+f(2)+f(4-1)=2+4+3=9

OpenStudy (anonymous):

how come f(3)=8

OpenStudy (sidsiddhartha):

it will be 10

OpenStudy (sidsiddhartha):

^*sorry

OpenStudy (anonymous):

yup i meant f(4-1)=f(3)=8

OpenStudy (anonymous):

how

OpenStudy (sidsiddhartha):

 ok??

OpenStudy (anonymous):

ok thanks now i got it so the answer of f(5)=22

OpenStudy (sidsiddhartha):

yup well done :)

OpenStudy (kirbykirby):

:)

OpenStudy (anonymous):

can you help me with another question please

OpenStudy (sidsiddhartha):

yeah u can ask it separately closing this one :)

OpenStudy (anonymous):

ok